Sheary, Inc., is proposing a rights offering. Presently, there are 400,000 shares outstanding at $54 each. There will be 25,000 new shares offered at $45 each. a. What is the new market value of the company? (Do not round intermediate calculations.) New market value $ b. How many rights are associated with one of the new shares? (Do not round intermediate calculations.) Number of rights needed rights c. What is the ex-rights price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Ex-rights price $ d. What is the value of a right? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.) Value of a right $
Answer #1

a. market value = shares*price = 400000*54=21600000

b. Rights = new shares/old shares = 25000/400000=0.0625

c.

ex rights price = (current MV+new shares*offer price)/(old shares+new shares)

=(21600000+25000*45)/(400000+25000)=53.47

d. Value of right = ((current price-subscription price)/( number of rights required to purchase one share

=((54-45)/(1/0.0625)=0.56